trace the curve y^2(1+x^2)=x^2(1-x^2) stating all the properties used
1
Expert's answer
2020-10-02T09:34:16-0400

"y^2 (1 + x^2) = x^2(1 - x^2)\\
\textsf{Differentiating both sides, we have;}\\
\begin{aligned}
2yy' + 2yy'x^2 + 2xy^2 &= 2x - 4x^3\\
y'(2y + 2yx^2) + 2xy^2 &= 2x - 4x^3
\end{aligned}\\

\textsf{At stationary point}\hspace{0.1cm} y' = 0.\\

\begin{aligned} 
y^2 &= \frac{2x - 4x^3}{2x}\\
y^2 &= 1 - 2x^2\\
x^2(1 - x^2) &= (1 - 2x^2)(1 + x^2)\\
x^2 - x^4 &= 1 + x^2 - 2x^2 - 2x^4\\
x^4 &= -2x^2 + 1\\
x^4 + 2x^2 - 1 &= 0
\end{aligned} \\

x^2 = \frac{-2 \pm \sqrt{8}}{2}\\
x^2 = -1 \pm \sqrt{2}\\

\textsf{The RHS of}\hspace{0.1cm} x^2 \textsf{cannot be negative}\\\textsf{in order to yield} \hspace{0.1cm}x\\
\therefore x = \pm\sqrt{\frac{-2 + \sqrt{8}}{2}} = \pm\sqrt{{-1 + \sqrt{2}}}\\

\therefore x = \pm\sqrt{\frac{-2 + \sqrt{8}}{2}} = \pm\sqrt{{-1 + \sqrt{2}}}\\


x = \pm \sqrt{{-1 + \sqrt{2}}} \approx \pm 0.6436\\




\textsf{At}\hspace{0.1cm} x = 0.6436, y = \pm 0.6436\sqrt{\frac{1 - 0.6436^2}{1 + 0.6436^2}}, y \approx \pm 0.4142.\\


\textsf{At}\hspace{0.1cm} x = -0.6436, y \approx \mp 0.4142 \\


\therefore \textsf{The stationary points are}\\(0.6436, 0.4142), (0.6436, -0.4142), \\(-0.6436, 0.4142),(-0.6436, -0.4142).\\


\textsf{Calculating the}\hspace{0.1cm}x \hspace{0.1cm} \textsf{and} \hspace{0.1cm}y\hspace{0.1cm}\textrm{intercepts}\\


\textsf{At}\hspace{0.1cm} x = 0, y = 0 \\



\textsf{At} \hspace{0.1cm}y = 0, x\sqrt{1 - x²} = 0 \Rightarrow x = 0, x = -1, 1. \\

\therefore \textsf{The}\hspace{0.1cm}x \hspace{0.1cm}\textrm{intercept}\hspace{0.1cm} \textsf{is}\hspace{0.1cm} (-1, 0),(0,0) \hspace{0.1cm}\& \hspace{0.1cm}(1, 0)\\ 


\textsf{and}\hspace{0.1cm}\textsf{the}\hspace{0.1cm} y\hspace{0.1cm}\textrm{intercept} \hspace{0.1cm} \textsf{is} \hspace{0.1cm} (0, 0). \\

\textsf{The graph is in the given below}"